# Resonant dynamics of spin cluster in a periodically driven one-dimensional Rydberg lattice

### v1

@@ IV.1 Resonance dynamics
To verify the above analysis and provide further conclusions, we perform numerical simulations of time evolutions under the effective Hamiltonian in Eq. (7). This allows us to handle a system with a larger $N$, and the results of spin cluster dynamics are expected to accord with that obtained from Hamiltonian in Eq. (1) under the condition of high frequency drive $\mathcal{V}_{0}\approx\omega\gg\mathit{\Omega}_{0}$.
In Fig. 2(a1), we present the result of time evolution of Rydberg density $\left<\hat{n}_{j}\right>(t)$ for the spin cluster that is initially prepared in a state of the form in Eq. (2) with fixed positions of domain walls. Here we take the interaction strength as $\mathcal{V}_{0}=5$, and then the resonant frequency is $\mathcal{F}=\sum_{l=1}^{r-1}\mathcal{V}_{0}/l^{6}\approx 5.0867$. The time evolution is numerically computed using the fourth-order Runge-Kutta method with time steps of length $\Delta t=2\times 10^{-4}$. At first glance, the profile of spin cluster experience ballistic expansion, while there is a change in Rydberg density at time $t\approx 13\mathit{\Omega}_{0}$, which is marked by a gray dashed line.
In order to gain a deeper understanding of the evolution of Rydberg density, we introduce the Rydberg density variance $\sigma(t)$ [44], which is defined as
@@
denotes the total Rydberg density. The density variance $\sigma(t)$ measures the spreading dynamics of spin cluster and relates to the mean square displacement [52, 53, 54, 55, 19], which is expected to increase over time as the form
@@
where $\sigma(0)$ is the density variance of the spin cluster at $t=0$. It is expected to observe $\beta=2$ for the ballistic transport. In Fig. 2(a2) and Fig. 3, we present the Rydberg density variance as a function of time on double-logarithmic scales and the total Rydberg density as a function of time, respectively, corresponding to the Rydberg density in Fig. 2(a1). We can see that before $t\approx 13\mathit{\Omega}_{0}$, the spin cluster did show ballistic expansion with $\beta=2.01$, while for a larger time $t\gtrsim 25\mathit{\Omega}_{0}$ the factor drops to $\beta=1.76$. Moreover, in Fig. 3, it is shown that the total Rydberg density is conserved before $t\approx 13\mathit{\Omega}_{0}$ and grows linearly for larger time. These behaviors arise because the number of spin clusters is conserved throughout the high-frequency drive process, and the size of the initial spin cluster $r_{0}$ is finite. Therefore when the size of the spin cluster shrinks to $r=1$, the effect of an effective hard-core repulsion potential appears. This is similar to the situation of the Rydberg lattice under the facilitation condition but without the driving field [31, 44]. Note that the hopping amplitude in Eq. (18) is half of the static case with $\omega=0$ for the hopping amplitude in Eq. (9). This leads to a time delay in the collision of the spin cluster’s domain walls.
@@
It is also worth noting that the hopping amplitude vanishes for the momentum component $k=\pi$, which would prevent any expansion or shrinkage of the spin cluster, leading to the ballistic expansion and the conservation of total Rydberg density of the spin cluster. In Figs. 2(b1)-(d1), Figs. 2(b2)-(d2) and Fig. 3, we present the numerical results of Rydberg density $\left<\hat{n}_{j}\right>(t)$, Rydberg density variance $\delta\sigma(t)$ and total Rydberg density $\mathcal{N}(t)$, respectively, for the initial Gaussian distribution in Eq. (6) with width $\sigma=4$ for different momentum $k_{0}$. The numerical results show that the cases with $k_{0}=0$ and $\pi/2$ are similar to that of the initial state $\left|\psi\left(j_{1,}j_{2}\right)\right$>: the spin clusters first experience ballistic expansion and then exhibit subdiffusive dynamics at a larger time. However, the results in Figs. 2(d1), (d2) and Fig. 3 indicate that for case with momentum $k_{0}=\pi$, the spin cluster indeed exhibit ballistic expansion, and the total Rydberg density is conserved at all times.
@@
Figure 4: Time evolution of Rydberg density $\left<\hat{n}_{j}\right$>, density variance $\delta\sigma(t)=\sigma(t)-\sigma(0)$ and autocorrelation function $\mathcal{A}(t)$ for the system near the resonant frequency of the driving field with different initial states. The frequency of the driving field is set as $\omega=5.1367$. The top panels (a1)-(d1) present time evolutions of Rydberg density $\left<\hat{n}_{j}\right$> for the initial excitations taken as the same as that in Fig. 2(a1)-(d1). The middle panels (a2)-(d2) show the density variance $\delta\sigma(t)$ corresponding to the Rydberg density in (a1)-(d1). The bottom panels show the autocorrelation function $\mathcal{A}(t)$ between the initial states and evolved states corresponding to the cases in (a1)-(d1). Other parameters are taken as $N=100$, $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$. The periodic boundary condition is taken for the system.
@@ IV.2 Dynamics near the resonant point
The results in Figs. 4(a1)-(a3) and Figs. 4(c1)-(c3) demonstrate that for both initial states of $\left|\psi\left(j_{1,}j_{2}\right)\right$> and Gaussian distribution with $k_{0}=\pi/2$, the systems exhibits the coexistence of oscillation and expansion dynamics of spin clusters. While for the Gaussian distribution with $k_{0}=0$ in Figs. 4(b1)-(b3), the system exhibit confinement dynamics of spin cluster without revival. These phenomena are mainly attributed to the collision of spin cluster domain walls. Notably, for the Gaussian distribution with $k_{0}=\pi$ in Figs. 4(d1)-(d3), the spin cluster exhibits oscillation dynamics with high amplitude revival, and as expected, the period is $T=2\pi\mathit{\Omega}_{0}/\delta\omega\approx 125.66\mathit{\Omega}_{0}$.
@@ V Summary and Discussion
In summary, we investigated the relaxation dynamics of the spin cluster in a 1D Rydberg lattice system under high-frequency driving field, and find a resonant frequency where the spin clusters exhibit ballistic expansion with half the spreading rate compared to the case under facilitation conditions before the domain walls of spin cluster collide. After the first collision, the spin clusters exhibit subdiffusive dynamics. While for an initial excitation with a Gaussian distribution of momentum $\pi$, the spin cluster exhibits ballistic expansion, and the total Rydberg density is conserved in the entire time domain. Moreover, we show that near the resonant driving frequency, the spin clusters manifest the coexistence of oscillation and expansion dynamics, and the Bloch-like oscillation of the spin cluster, for different initial excitations. These results demonstrate the rich dynamic behaviors in the driven Rydberg spin lattices.
Our conclusions are applicable to the 1D Rydberg lattice with other types of long-range interactions, such as the dipole-dipole interaction with the form $V_{i,j}=\mathcal{V}_{0}/\left|i-j\right|^{3}$. The resonant dynamics of the spin cluster can also be observed in the transverse Ising chain in the presence of a longitudinal field [56, 57, 58, 59, 60, 61, 62]. In this case, the longitudinal field can be periodically modulated to be resonant with the transverse field. In the future, it would be interesting to study the impacts of spin-phonon interactions [44, 49] on the resonant dynamics in the driven Rydberg lattices.
@@ Acknowledgements.
This work was supported by the Natural Science Foundation of Guangdong Province through Grant Nos. 2024A1515110222 and 2023A1515010770, the National Natural Science Foundation of China (NSFC) through Grant Nos. 12305013 and 12274077, the Research Fund of Guangdong-HongKong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology (No. 2020B1212030010).
APPENDIX The transition probabilities
@@
While for Hamiltonian $\hat{H}_{j}(t)=\mathit{\Omega}_{0}\cos\left(\omega t\right)\hat{\sigma}_{j}^{x}/2+\mathit{\Delta}_{\mathrm{eff}}\hat{n}_{j}$, the transition probability can be evaluated by the time-dependent perturbation theory [63]. Under the condition $\mathit{\Omega}_{0}/2\ll\mathit{\Delta}_{\mathrm{eff}}$, the first term of $\hat{H}_{j}(t)$ can be treated as a perturbation. Up to first order, the amplitude of transition from the Rydberg state to ground state is

### v2

@@ IV.1 Resonance dynamics
To verify the above analysis and provide further conclusions, we perform numerical simulations of time evolutions under the effective Hamiltonian in Eq. (7). This allows us to handle a system with a larger $N$, and the results of spin cluster dynamics are expected to accord with that obtained from the Hamiltonian in Eq. (1) under the condition of high-frequency drive $\mathcal{V}_{0}\approx\omega\gg\mathit{\Omega}_{0}$.
In Fig. 3(a1), we present the result of time evolution of Rydberg density $\left<\hat{n}_{j}\right>(t)$ for the spin cluster that is initially prepared in a state of the form in Eq. (2) with fixed positions of domain walls. Here we take the interaction strength as $\mathcal{V}_{0}=5$, and then the resonant frequency is $\mathcal{F}=\sum_{l=1}^{r-1}\mathcal{V}_{0}/l^{6}\approx 5.0867$. The time evolution is numerically computed using the fourth-order Runge-Kutta method with time steps of length $\Delta t=2\times 10^{-4}$. At first glance, the profile of the spin cluster experiences ballistic expansion, while there is a change in Rydberg density at time $t\approx 13\mathit{\Omega}_{0}$, which is marked by a gray dashed line.
In order to gain a deeper understanding of the evolution of Rydberg density, we introduce the Rydberg density variance $\sigma(t)$ [Magoni2024], which is defined as
@@
denotes the total Rydberg density. The density variance $\sigma(t)$ measures the spreading dynamics of spin cluster and relates to the mean square displacement [steinigeweg2017real, Lev2017, lm2022logarithmic, sierant2023slow, zhang2024two], which is expected to increase over time as the form
@@
where $\sigma(0)$ is the density variance of the spin cluster at $t=0$. It is expected to observe $\beta=2$ for the ballistic transport. In Fig. 3(a2) and Fig. 4, we present the Rydberg density variance as a function of time on double-logarithmic scales and the total Rydberg density as a function of time, respectively, corresponding to the Rydberg density in Fig. 3(a1). We can see that before $t\approx 13\mathit{\Omega}_{0}$, the spin cluster did show ballistic expansion with $\beta=2.01$, while for a larger time $t\gtrsim 25\mathit{\Omega}_{0}$ the factor drops to $\beta=1.76$. Moreover, in Fig. 4, it is shown that the total Rydberg density is conserved before $t\approx 13\mathit{\Omega}_{0}$ and grows linearly for larger time. These behaviors arise because the number of spin clusters is conserved throughout the high-frequency drive process, and the size of the initial spin cluster $r_{0}$ is finite. Therefore when the size of the spin cluster shrinks to $r=1$, the effect of an effective hard-core repulsion potential appears. This is similar to the situation of the Rydberg lattice under the facilitation condition but without the driving field [Marcuzzi2017, Magoni2024]. Note that the hopping amplitude in Eq. (20) is half of the static case with $\omega=0$ for the hopping amplitude in Eq. (11). This leads to a time delay in the collision of the spin cluster’s domain walls.
@@
It is also worth noting that the hopping amplitude vanishes for the momentum component $k=\pi$, which would prevent any expansion or shrinkage of the spin cluster, leading to the ballistic expansion and the conservation of total Rydberg density of the spin cluster. In Figs. 3(b1)-(d1), Figs. 3(b2)-(d2) and Fig. 4, we present the numerical results of Rydberg density $\left<\hat{n}_{j}\right>(t)$, Rydberg density variance $\delta\sigma(t)$ and total Rydberg density $\mathcal{N}(t)$, respectively, for the initial Gaussian distribution in Eq. (6) with width $d=4$ for different momentum $k_{0}$. The numerical results show that the cases with $k_{0}=0$ and $\pi/2$ are similar to that of the initial state $\left|\psi\left(j_{1,}j_{2}\right)\right$>: the spin clusters first experience ballistic expansion and then exhibit subdiffusive dynamics at a larger time. However, the results in Figs. 3(d1), (d2) and Fig. 4 indicate that for case with momentum $k_{0}=\pi$, the spin cluster indeed exhibit ballistic expansion, and the total Rydberg density is conserved at all times. The primary interest in this effect stems from its occurrence in a system featuring the combined effects of long-range interactions and time-dependent driving.
@@
Figure 6: Time evolution of Rydberg density $\left<\hat{n}_{j}\right$>, density variance $\delta\sigma(t)=\sigma(t)-\sigma(0)$ for the system near the resonant frequency of the driving field with spin cluster initial states $\left|\psi\left(j_{1},j_{2}\right)\right$>. The frequency of the driving field is set as $\omega=5.1367$. The top panels (a1) and (b1) present time evolutions of Rydberg density $\left<\hat{n}_{j}\right$> for the initial excitations with $(j_{1},j_{2})=(41,60)$ and $(j_{1},j_{2})=(31,70)$, respectively. The bottom panels (a2) and (b2) show the density variance $\delta\sigma(t)$ corresponding to the Rydberg density in (a1) and (b1). The size of the system is set as $N=100$, the strength of the field and interaction are taken as $\mathit{\Omega}_{0}=1$ and $\mathcal{V}_{0}=5$, respectively, and the periodic boundary condition is taken for the system. These indicate that the domain walls of spin cluster can no longer collide when the initial spin cluster size $r_{0}=j_{2}-j_{1}+1$ is larger than the amplitude of oscillation.
@@ IV.2 Dynamics near the resonant point
The results in Figs. 5(a1)-(a3) and Figs. 5(c1)-(c3) demonstrate that for both initial states of $\left|\psi\left(j_{1,}j_{2}\right)\right$> and Gaussian distribution with $k_{0}=\pi/2$, the systems exhibits the coexistence of oscillation and expansion dynamics of spin clusters, while for the Gaussian distribution with $k_{0}=0$ in Figs. 5(b1)-(b3), the systems exhibit confinement dynamics of spin cluster without revival. These phenomena are mainly attributed to the collision of spin cluster domain walls. Notably, for the Gaussian distribution with $k_{0}=\pi$ in Figs. 5(d1)-(d3), the spin cluster exhibits oscillation dynamics with high amplitude revival, and as expected, the period is $T=2\pi\mathit{\Omega}_{0}/\delta\omega\approx 125.66\mathit{\Omega}_{0}$.
Moreover, the initial spin cluster size $r_{0}$ is closely related to the collision of domain walls of spin cluster. For the resonant case, the collision time of domain walls is proportional to $r_{0}$, while for oscillation dynamics near the resonant frequency, the domain walls can no longer collide when $r_{0}$ is larger than the amplitude of oscillation. In this case, the spin cluster does not expand and only oscillation dynamics exists. In Fig. 6, we present the numerical results for time evolution of Rydberg density and density variance near the resonant frequency for $r_{0}=20$ and $r_{0}=40$. In comparison with the cases with $r_{0}=10$ in Fig. 5(a1) and $r_{0}=20$ in Fig. 6(a1), we can see that for the case with $r_{0}=40$, the two domain walls move independently in Figs. 6(b1) and (b2), and the oscillation amplitude of density variance $\delta\sigma(t)$ does not increase as time.
@@ V Summary and Discussion
In summary, we investigate the relaxation dynamics of the spin cluster in a 1D Rydberg lattice system under high-frequency driving field, and find a resonant frequency where the spin clusters exhibit ballistic expansion with half the spreading rate compared to the case under facilitation conditions before the domain walls of the spin cluster collide. After the first collision, the spin clusters exhibit subdiffusive dynamics, while for an initial excitation with a Gaussian distribution of momentum $\pi$, the spin cluster exhibits ballistic expansion, and the total Rydberg density is conserved in the entire time domain. Moreover, we show that near the resonant driving frequency, the spin clusters manifest the coexistence of oscillation and expansion dynamics, and the Bloch-like oscillation of the spin cluster, for different initial excitations.
These features are highly promising for experimental verification in future research. In the Rydberg lattice system, the time-dependent driving fields can be readily implemented experimentally through the Rydberg lasers with time-varying Rabi frequency [Bluvstein2021]. The single-site addressability allows for the individual excitation of atoms from the ground to the Rydberg state [Bernien2017, de2017optical], and thus the spin cluster state is feasible in experiments. While the preparation of initial state with Gaussian distribution is more difficult, the former work [surace2021scattering] suggests a protocol to generating a narrow momentum distribution of spin wave packet. For example, on the left side of the spin chain, a spatially localized spin flip is generated by spatially inhomogeneous fields. The sharp variation in the field parameters allows only components with specific momenta to propagate to the right, thereby achieving momentum filtering. It is also possible to implement a similar protocol to our work by introducing an inhomogeneity into the lattice.
VI Conclusion
The modulated drive in our work does not require modulating the detuning precisely to control the dynamics, and may provide another option or degree of freedom in parameter for manipulating the dynamics of spin clusters in the 1D Rydberg lattice. Moreover, we find that there is a time delay in the collision of the spin cluster’s domain walls in comparison to the case with the facilitation condition because the effective hopping amplitude is half of that of the static case. it is noteworthy that ballistic transport of spin clusters occurs in a system with long-range interactions and time-dependent driving. These results demonstrate the rich dynamic behaviors in the driven Rydberg spin lattices.
Our conclusions are applicable to the 1D Rydberg lattice with other types of long-range interactions, such as the dipole-dipole interaction with the form $V_{i,j}=\mathcal{V}_{0}/\left|i-j\right|^{3}$. The resonant dynamics of the spin cluster can also be observed in the transverse Ising chain in the presence of a longitudinal field [simon2011quantum, cai2011, kormos2017real, verdel2020, lerose2020quasilocalized, verdel2023, lagnese2024detecting]. In this case, the longitudinal field can be periodically modulated to be resonant with the transverse field. In the future, it would be interesting to study the impacts of spin-phonon interactions [Magoni2024, Brady2025] on the resonant dynamics in the driven Rydberg lattices.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12505015, No. 12305013, and No. 12274077), the Natural Science Foundation of Guangdong Province (Grants No. 2024A1515110222, and No. 2023A1515010770), and the Research Fund of Guangdong-HongKong-Macao Joint Laboratory for Intelligent Micro-Nano Optoelectronic Technology (No. 2020B1212030010).
APPENDIX A The transition probabilities
@@
However, for Hamiltonian $\hat{H}_{j}(t)=\mathit{\Omega}_{0}\cos\left(\omega t\right)\hat{\sigma}_{j}^{x}/2+\mathit{\Delta}_{\mathrm{eff}}\hat{n}_{j}$, the transition probability can be evaluated by the time-dependent perturbation theory [sakurai2020modern]. Under the condition $\mathit{\Omega}_{0}/2\ll\mathit{\Delta}_{\mathrm{eff}}$, the first term of $\hat{H}_{j}(t)$ can be treated as a perturbation. Up to first order, the amplitude of transition from the Rydberg state to ground state is
@@
APPENDIX B Details of numerical simulations
Here we present the details of the numerical simulations of time evolution with time-dependent Hamiltonian $\hat{\mathcal{H}}\left(t\right)$. The evolved state at time $t+\Delta t$ can be computed by the standard fourth-order Runge–Kutta method with the expansion [carpenter1994fourth]
$|\Psi\left(t+\Delta t\right)\rangle=$ $|\Psi\left(t\right)\rangle+\frac{1}{6}|\Psi^{\left(1\right)}\rangle+\frac{1}{3}|\Psi^{\left(2\right)}\rangle$ (B1)
$+\frac{1}{3}|\Psi^{\left(3\right)}\rangle+\frac{1}{6}|\Psi^{\left(4\right)}\rangle+O(\Delta t^{5}),$
where four auxiliary vectors are
$\begin{aligned} |\Psi^{\left(1\right)}\rangle&=-\mathrm{i}\hat{\mathcal{H}}\left(t\right)|\Psi\left(t\right)\rangle\Delta t\\ |\Psi^{\left(2\right)}\rangle&=-\mathrm{i}\hat{\mathcal{H}}\left(t+\frac{\Delta t}{2}\right)\left(|\Psi\left(t\right)\rangle+\frac{1}{2}|\Psi^{\left(1\right)}\rangle\right)\Delta t\\ |\Psi^{\left(3\right)}\rangle&=-\mathrm{i}\hat{\mathcal{H}}\left(t+\frac{\Delta t}{2}\right)\left(|\Psi\left(t\right)\rangle+\frac{1}{2}|\Psi^{\left(2\right)}\rangle\right)\Delta t\\ |\Psi^{\left(4\right)}\rangle&=-\mathrm{i}\hat{\mathcal{H}}\left(t+\Delta t\right)\left(|\Psi\left(t\right)\rangle+|\Psi^{\left(3\right)}\rangle\right)\Delta t\end{aligned}.$ (B2)
The time steps of length is taken as $\Delta t=4\times 10^{-4}$ for the numerical computations. In the implementation of this method, one can exploit the sparsity of the matrix representation of the Hamiltonian to save the computer memory and time.